OpenStudy (anonymous):

trigonometric identities: cos^4a-sen^4a+1=2cos^2a

OpenStudy (jamesj):

You mean sin Here's a hint. Write c = cos a, and s = sin a. Then c^4 - s^4 = (c^2 - s^2)(c^2 + s^2) = c^2 - s^2

OpenStudy (anonymous):

thanks, but i already had that part :S i needed the rest, im on c^2-s^2=2c^2